OpenStudy (anonymous):

Rewrite sin(cos^-1x+sin^-1x) as an algebraic expression.

OpenStudy (anonymous):

1 according to Mathematica.

OpenStudy (anonymous):

how do they solve it? I am having trouble on where to begin...

OpenStudy (freckles):

use sum identity for sine

OpenStudy (anonymous):

\[\sin^{-1} x+\cos^{-1} x=\frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

\[\sin \frac{ \pi }{ 2 }=1\]

OpenStudy (anonymous):

\[put~\sin^{-1} x=t,x=\sin t=\cos \left( \frac{ \pi }{ 2 }-t \right)\] \[or~\cos^{-1} x=\frac{ \pi }{ 2 }-t=\frac{ \pi }{ 2 }-\sin^{-1} x\] \[or~\sin^{-1} x+\cos^{-1} x=\frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

Thanks , but I figured it out.

OpenStudy (anonymous):

yw